### v2

Abstract
Understanding firm conduct is crucial for industrial organization and antitrust policy. In this article, we develop a testing procedure based on the Rivers and Vuong non-nested model selection framework. Unlike existing methods that require estimating the demand and supply system, our approach compares the model fit of two first-stage price regressions. Through an extensive Monte Carlo study, we demonstrate that our test performs comparably to, or outperforms, existing methods in detecting collusion across various collusive scenarios. The results are robust to model misspecification, alternative functional forms for instruments, and data limitations. By simplifying the diagnosis of firm behavior, our method provides an efficient tool for researchers and regulators to assess industry conduct. Additionally, our approach offers a practical guideline for enhancing the strength of BLP-style instruments in demand estimation: once collusion is detected, researchers are advised to incorporate the product characteristics of colluding partners into own-firm instruments while excluding them from other-firm instruments.
Keywords: firm conduct, BLP-style instruments, Rivers and Vuong test, discrete choice models, demand estimation
JEL Classification Numbers: C36, L13, L21, L41
This version: March 1st 2025
First version: January 8th 2025
1 Introduction
Understanding industry conduct – that is, the extent to which firms collude in a goods market – is one of the central issues in the study of industrial organization. This issue also carries practical importance, as antitrust policies and regulations based on incorrect conclusions about industry conduct could harm consumers and reduce market efficiency. For this reason, IO researchers and regulatory authorities have shown significant interest in accurately evaluating firm conduct. In response to this interest, various approaches to testing firm conduct have been developed.
Since existing tests on industry conduct (e.g., duarte2024testing, backus2021common, dearing2024learning) require estimating the demand and supply system – a task that remains challenging despite advances in modeling and computing over the past three decades – their testing power can be reduced by model misspecification. Furthermore, the true firm conduct is often unknown, complicating the proper construction of BLP-style instruments – functions of exogenous product characteristics, such as the sums of characteristics of products within and across firms. For instance, if a researcher estimates demand in a differentiated product market and assumes that firms engage in price competition when they actually collude, the BLP-style instruments may not be constructed effectively, resulting in reduced identification power.
In this article, we propose a practical and powerful method for testing firm conduct that is also useful for designing more effective BLP-style instruments. The intuition behind our testing procedure is that if firms engage in price competition, the price effect of a change in own-firm product characteristics is not only larger in magnitude but also opposite in direction to the price effect of a change in other-firm product characteristics (Berry95; gandhi2021empirical). In contrast, in the case of collusion, the two price effects are similar in magnitude (assuming the products are equally substitutable) and share the same direction. Therefore, treating the product attributes of other firms as own attributes when constructing BLP-style instruments would enhance their strength in the first-stage price regression in cases of collusion. On the other hand, it would weaken the instruments’ strength if the firms engage in price competition. Consequently, the strength of BLP-style instruments hinges on whether the assumption about firm conduct – unknown to the researcher – is correctly imposed. In other words, we can infer whether firms compete or collude by comparing the effectiveness of instruments constructed under alternative firm conduct scenarios.
Based on the above idea, we develop a testing procedure by applying the non-nested model selection method proposed by vuong1989likelihood and rivers2002model (the RV test). The first step in our procedure is to construct two BLP-style instruments: one incorporating only own-firm product attributes (competition IVs) and the other incorporating both the characteristics of own products and those of suspected colluding partners (collusion IVs). The next step is to run two linear price regressions separately using each of the two instrument sets. Finally, the last stage involves constructing the RV test statistic based on the objective function values of the two regression models. A statistically significant positive (negative) value can be interpreted as evidence of collusive (competitive) behavior among the suspected firms. The primary advantage of our method over existing approaches built upon equilibrium markup conditions is that it does not require estimating the demand and supply system, as it simply compares the model fit of the two first-stage price regressions.
To evaluate the finite-sample performance of our proposed test statistic, we conduct an extensive Monte Carlo study. More specifically, we consider an indirect utility function with a normally distributed random coefficient for the price (as well as a simple logit demand model) and a linear marginal cost function. We then simulate 500 datasets for various Monte Carlo configurations, each representing a unique market condition (in terms of the degree of market concentration, the share of colluding firms among all firms, the extent to which a firm internalizes its colluding partners’ profits). Using each simulated dataset, we calculate both our test statistic and the existing statistic. Finally, we compute the median values of the test statistics for each Monte Carlo configuration and examine how they change as industry conduct varies.
The results, which are robust to model misspecification, data limitations (e.g., the availability of only a few markets), and the use of alternative functional forms for BLP-style instruments, reveal that our testing procedure performs comparably to, or outperforms, existing approaches in detecting collusion under various market conditions. The high testing power of our method, combined with its ease of implementation – requiring only a comparison of the model fit of two first-stage price regressions without estimating the demand and supply system – and minimal data requirements, makes it a practical tool for the preliminary diagnosis of firm conduct. Once our test statistic indicates the presence of collusion among suspected firms in the market, researchers and regulatory authorities can proceed with existing approaches to determine which specific conduct model best fits the observed data. In this way, the two methods complement each other and collectively create a more practical, efficient, and powerful framework for conduct testing.
Our Monte Carlo study also reveals that, as expected, the use of collusion IVs instead of competition IVs improves the estimation performance of demand parameters under various Bertrand-Nash collusive scenarios. This finding suggests that researchers can enhance the strength of BLP-style instruments by designing them to accurately reflect actual industry conduct. For instance, after finding suggestive evidence of collusion using our testing procedure, researchers are advised to incorporate the product characteristics of colluding partners into own-firm instruments while excluding them from other-firm instruments.
We apply our testing method to study firm conduct in two differentiated product markets in South Korea: the new passenger car market and the instant noodles market. Our test rejects the hypotheses of brand-level profit maximization and price collusion in favor of joint profit maximization at the parent company level in the car market. In addition, our results suggest that, despite suspicions from the Korean Fair Trade Commission, instant noodle manufacturers have not engaged in any price collusion. Our findings align with anecdotal evidence of firm conduct in each market, court rulings, and previous empirical findings that instant noodle prices remain significantly below the collusive level (kim2024collusion).
This article proceeds as follows. In the next section, we review previous literature and delineate the contributions of our study. In Section 3, we provide a conceptual framework for our testing procedure which we formally propose in Section 4. Based on an extensive Monte Carlo study whose setup is elaborated in Section 5, we examine the performance of our testing method and compare it with that of existing procedures in Section 6. We then conduct another simulation study in Section 7 to provide further intuition behind our approach. In Section 8, we apply our method to study industry conduct in two differentiated product markets. Finally, we conclude in Section 9.
2 Related literature
This article is closely related to the empirical literature on implementing the RV approach for testing firm conduct. [^1] As discussed in duarte2024testing, alternative testing procedures include (i) the Estimation-Based (EB) test (pakes2017empirical), (ii) the Anderson-Rubin (AR) test (anderson1949estimation), and (iii) the Cox test (cox1961tests; 98610483-571d-3668-af2c-db6b7ff86a65). Examples of empirical applications of these approaches include miller2017understanding (EB test), bergquist2020competition (AR test), and Villas-Boas07 (Cox test). An advantage of the RV test over these methods is that inference on conduct remains robust even under misspecification. bonnet2010inference investigated vertical relationships between manufacturers and retailers in the French bottled water market, concluding that manufacturers use two-part tariff contracts with resale price maintenance. In the premium ice cream market, sullivan2017ice found evidence that Ben & Jerry’s and Häagen-Dazs exhibited behaviors consistent with full coordination on both pricing and product choice decisions. [^2] To examine coordinated behavior in product choice decisions, sullivan2017ice modeled endogenous product choices and estimated bounds for the fixed cost parameter using the partial identification approaches employed in eizenberg2014upstream and wollmann2018trucks. He then implemented the RV test extended by shi2015nondegenerate within a moment inequalities framework. hu2014collusion examined price collusion within or across large corporate groups in the Chinese passenger-vehicle industry and found no evidence of collusive pricing behavior. These previous studies backed out marginal costs from demand estimates under each assumed firm conduct model, constructed moment-based objective functions, and determined which model provided a better fit.
Recent studies on the RV test approach have built upon the falsifiable restrictions proposed by berry2014identification to distinguish between alternative conduct models in differentiated goods markets. These studies have further examined the identification conditions required for testing, the selection of appropriate instruments, and the econometric properties of the RV approach. dearing2024learning identified cost pass-through as a key determinant for selecting instruments to falsify models of firm behavior. duarte2024testing analyzed the properties of the RV test statistic under weak instrument conditions and developed a robust inference framework for the RV approach in such scenarios. They also highlighted the advantages of the RV approach under model misspecification compared to other testing methods. Furthermore, backus2021common refined the RV testing procedure by incorporating flexible nonparametric functional forms for marginal costs and proposing instrument functional forms that utilize scalar moment values to enhance the testing power of the RV framework. Our approach, while relying solely on first-stage price regressions, performs comparably to these methods, which are built upon equilibrium markup conditions, and thus offers a simple yet effective diagnostic for detecting collusive pricing behavior.
This article also contributes to the broader empirical industrial organization (IO) literature on assessing firm conduct in differentiated product markets, a longstanding and central topic of inquiry in the field and among antitrust authorities. Examples include airlines (ciliberto2014does), automobiles (bresnahan1987competition; verboven1996international; sudhir2001competitive; hu2014collusion), beer (5c9e3160-6f4b-35ab-b813-4b5f985f6fe0; slade2004; rojas2008price; miller2017understanding; miller2021oligopolistic), instant noodles (kim2024collusion), ready-to-eat (RTE) cereal (Nevo01; backus2021common; michel2024estimating), soft drinks (gasmi1992econometric), and ice cream (sullivan2017ice). These studies relied on structural models of demand and supply to identify the firm behavior that best explains the observed market outcomes. In contrast, our proposed method requires only a pair of price regressions, making it straightforward to implement.
Finally, literature on demand estimation with aggregate data in the discrete choice framework also bears on this article. Prior works (e.g., berry1999voluntary, berry2014identification, reynaert2014improving, armstrong2016large, gandhi2019measuring, and conlon2020best) have addressed challenges arising from weak instruments in non-linear GMM estimation. Our work contributes to this field by enhancing the understanding of instruments and improving estimation performance, demonstrating that the strength of BLP-style instruments depends on the correctness of the imposed assumption on industry conduct.
3 Conceptual framework
3.1 Discrete choice in differentiated product markets
There are $J_{t}+1$ differentiated products in market $t=1,2,\ldots,T$. For product $j=0,1,\ldots,J_{t}$, there are observed attributes $(x_{jt},p_{jt})$ and an unobserved component $\xi_{jt}$. The observed attributes are grouped into two parts: (i) the exogenous part, $x_{jt}$, which is uncorrelated with the unobservables $\bm{\xi}_{t}=(\xi_{0t},\xi_{1t},\ldots,\xi_{Jt})$ and (ii) the endogenous price, $p_{jt}$, which is correlated with $\bm{\xi}_{t}$. For each product, there are $R$ exogenous attributes in $x_{jt}=(x_{jt}^{(1)},\ldots,x_{jt}^{(r)},\ldots,x_{jt}^{(R)})$. Market $t$ is defined by the set of products offered, denoted by $\{0\}\cup\mathscr{J}_{t}$, the set of attributes of these products, denoted by $\chi_{t}=\{(x_{0t},p_{0t},\xi_{0t}),\ldots,(x_{Jt},p_{Jt},\xi_{Jt})\}$, and the set of characteristics of consumers, denoted by $V_{t}$.
A consumer $i$ with characteristics $v_{i}\in V_{t}$ has the indirect utility from consuming product $j$ given by $u_{ijt}=U(x_{jt},p_{jt},\xi_{jt},v_{i})$. A consumer purchases the product that maximizes his or her utility. The predicted market share of product $j$, denoted by $s_{jt}$, is the fraction of market consumers who prefer good $j$ over all other $J_{t}$ products:
$s_{jt}=\Pr(v_{i}\in V_{t}:U(x_{jt},p_{jt},\xi_{jt},v_{i})\geq U(x_{kt},p_{kt},\xi_{kt},v_{i})\quad\forall k\in\{0\}\cup\mathscr{J}_{t})$
Consumer characteristics, $v_{i}$, are decomposed into a common component shared across all individuals and heterogeneous components: (i) random utility components, $\nu_{i}=(\nu_{ip},\nu_{i1},\ldots,\nu_{iR})$, capturing heterogeneous valuations of $(p_{jt},x_{jt})$; and (ii) an idiosyncratic taste shock, $\varepsilon_{ijt}$. The most commonly used utility specification in the literature is given by:
$u_{ijt}=\delta_{jt}+\nu_{ip}p_{jt}+\sum\limits_{r}\nu_{ir}x_{jt}^{(r)}+\varepsilon_{ijt},$ (1)
where $\delta_{jt}=x_{jt}\beta-\alpha p_{jt}+\xi_{jt}$ is the mean utility of the product, which also absorbs the common component of $v_{i}$. Since the heterogeneous components $\nu_{i}$ and $\varepsilon_{ijt}$ are not directly observed in aggregated market-level data, we assume their distributions. Let $\nu_{i}$ follow a mean-zero multivariate normal distribution, [^3] Demographic-specific valuations of $(x_{jt},p_{jt})$ can also be modeled either by drawing from an empirical distribution or by imposing distributional assumptions on the demographics of market $t$. while $\varepsilon_{ijt}$ are i.i.d. (across both consumers and products) extreme value distributed.
We normalize the indirect utility from consuming product $j=0$ by setting $u_{i0t}=\varepsilon_{i0t}$ for all consumers and refer to it as the outside option. Assuming that the two random components, $\nu$ and $\varepsilon$, are independent, the predicted market share of product $j$ is given by
[…]
where $F(\nu_{i};\theta_{2})$ is the joint cumulative distribution function of $\nu_{i}$, governed by a vector of non-linear parameters $\theta_{2}$. Bold-faced letters indicate that the predicted market share depends on the vectors of mean utilities and attributes of all products in market $t$.
Berry94 demonstrated that, for a given non-linear parameter vector $\theta_{2}$, there exists a unique vector of mean utilities $\bm{\delta}_{t}$ that equates the observed market shares with the predicted market shares. Specifically, let $\mathbf{S}_{t}=(S_{0t},S_{1t},\ldots,S_{Jt})$ denote the vector of observed market shares of products in market $t$. Then, $\mathbf{s}_{t}(\bm{\delta}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})=\mathbf{S}_{t}$. This relationship allows us to invert the demand system in market $t$ and solve for $\delta_{jt}(\mathbf{S}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})=x_{jt}\beta-\alpha p_{jt}+\xi_{jt}$ for all $j\in\mathscr{J}_{t}$. [^4] This invertibility result also holds for more general specifications of $U(\mathbf{x}_{t},\mathbf{p}_{t},\bm{\xi}_{t},v_{i})$. Specifically, when the utility specification induces connected substitutes relationships among products in the market, invertibility is ensured. See berry2013connected for further details. Finally, for a given parameter vector $\theta=(\theta_{1},\theta_{2})$, where $\theta_{1}=(\alpha,\beta)$ represents the linear parameters, we compute $\xi_{jt}(\theta)=\delta_{jt}(\mathbf{S}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})-(x_{jt}\beta-\alpha p_{jt})$.
Utility parameters are estimated using a GMM framework based on the moment restrictions at the true parameter $\theta=\theta_{0}$ given by:
$E[\xi_{jt}(\theta_{0})|\mathbf{Z}_{jt}]=0,$
where $\mathbf{Z}_{jt}=(x_{jt},\mathbf{z}_{jt})$ is a vector of instruments, and $\mathbf{z}_{jt}$ is a vector of excluded instruments. For the price of product $p_{jt}$, Berry95 used the sum of attributes across own-firm products and the sum of attributes across other-firm products as excluded instruments. [^5] Importantly, these instruments also help address endogeneity arising from observed market shares and prices of other products. Specifically, $\xi_{jt}(\theta)$ is a function of $\mathbf{S}_{t}$ and $\mathbf{p}_{t}$. Therefore, estimating the inverted demand system with a flexible utility specification requires instruments that are exogenous to these variables as well. See berry2014identification for identification results and instrument requirements in nonparametric settings. Specifically, $\mathbf{z}_{jt}=(\mathbf{z}^{own}_{jt},\mathbf{z}^{other}_{jt})$ is given by
$\mathbf{z}_{jt}^{own}=\left(\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}x_{kt}^{(1)},\ldots,\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}x_{kt}^{(R)}\right),\quad\mathbf{z}_{jt}^{other}=\left(\sum_{k\in\mathscr{J}_{t}\setminus\mathscr{J}_{ft}}x_{kt}^{(1)},\ldots,\sum_{k\in\mathscr{J}_{t}\setminus\mathscr{J}_{ft}}x_{kt}^{(R)}\right),$ (3)
where $\mathscr{J}_{ft}\setminus\{j\}$ denotes the set of products, excluding $j$, offered by firm $f$ in market $t$. These instruments, constructed as functions of all product attributes in the market, are referred to as BLP-style instruments throughout this article.
3.2 Oligopoly model
There are $F_{t}$ firms in market $t$, which has a size of $M_{t}$, measuring the number of consumers. Importantly, firms may engage in collusion, internalizing the profits of their colluding partners, while the true industry conduct is unknown to the researcher or regulatory authority. The variable profit of firm $f=1,2,\ldots,F_{t}$ is given by:
$\Pi_{ft}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t},\phi)$ $=\sum_{k\in\mathscr{J}_{ft}}(p_{k}-mc_{k})\cdot s_{kt}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t})\cdot M_{t}$ (4)
$+\phi\cdot\sum_{k\in\mathscr{J}_{f_{c}t}\setminus\mathscr{J}_{ft}}(p_{k}-mc_{k})\cdot s_{kt}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t})\cdot M_{t},$
where $f_{c}$ is a firm index that treats colluding entities as a single entity. For instance, if firms 1 and 2 collude, then $f_{c}=1$ for $f=1$ and $f=2$, resulting in $f_{c}=1,3,\ldots,F_{t}$.
The degree of profit internalization is measured by $\phi\in[0,1]$. As shown in equation (4), firm $f$ partially internalizes the profits of its colluding partners, weighting them by $\phi$ as part of its own profit. Firms in market $t$ set profit-maximizing prices simultaneously. The Bertrand-Nash equilibrium prices in market $t$, and consequently, the vector of equilibrium markups, are determined by the following First-Order Conditions (FOCs):
$\mathbf{p}_{t}-\mathbf{mc}_{t}=-\left[\mathscr{H}_{t}\odot\frac{\partial\mathbf{s}_{t}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t})}{\partial\mathbf{p}_{t}}\right]^{-1}\mathbf{s}_{t}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t}),$ (5)
where $\mathbf{mc}_{t}$ is a vector of marginal costs, $\odot$ denotes component-wise multiplication, and $\mathscr{H}_{t}$ is an unknown (to the researcher) $J_{t}\times J_{t}$ ownership matrix. For example, $\mathscr{H}_{t,jk}=1$ if $j,k\in\mathscr{J}_{ft}$ and 00 otherwise in the case of competition ($\phi=0$), while $\mathscr{H}_{t}$ is a matrix of ones in the case of full collusion ($\phi=1$ and $f_{c}=1$ for all $f$). As shown in the markup equation (5) above, the endogeneity of prices arises from the fact that firms set these prices based on demand, making them functions of the unobservable components $\bm{\xi}$.
The markup equation (5) also demonstrates that in an oligopoly, pricing is influenced by the proximity of a product to its substitutes within the product characteristics space. Products facing close competition tend to have lower markups and prices, while those that are significantly differentiated can command higher markups and prices. More importantly, the equation highlights that the impact of these characteristics on pricing depends on the unknown ownership structure.
3.3 Firm conduct and the strength of BLP-style instruments
Logit case
To clarify this point, assume a simple logit model where $v_{ip}=v_{ir}=0$ in the utility specification (1). Under this assumption, the equilibrium markups for firm $f$ in market $t$ are derived as:
$\mathbf{p}_{ft}-\mathbf{mc}_{ft}=\left\{\begin{array}[]{ll}\frac{1}{\alpha}\bigg{(}\frac{1+\sum_{k\in\text{market}}X_{kt}}{1+\sum_{k\in\text{market}}X_{kt}-\sum_{k\in\text{own}}X_{kt}}\bigg{)}\cdot\mathbf{1}_{ft}&\textrm{under competition}\vspace{0.1in}\\ \frac{1}{\alpha}\bigg{(}\frac{1+\sum_{k\in\text{market}}X_{kt}}{1+\sum_{k\in\text{market}}X_{kt}-\big{(}\sum_{k\in\text{own}}X_{kt}+\sum_{k\in\text{colluding}}X_{kt}\big{)}}\bigg{)}\cdot\mathbf{1}_{ft}&\textrm{under full collusion}\\ \end{array}\right.$ (8)
where $\mathbf{1}_{ft}$ is a $J_{ft}\times 1$ unit vector and $X_{kt}=\exp(\delta_{kt})$. Note that $\delta_{kt}$ is a function of the attributes of product $k$. One can see that the BLP-style instruments, $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$, affect the equilibrium markups and consequently, the equilibrium prices. Moreover, since the equilibrium markup is a function of $\mathscr{H}_{t}$, the identifying power of BLP-style instruments depends not only on the true market structure and competition level, but also on whether the industry conduct assumption upon which the BLP-style instruments are constructed is correct. [^6] It also depends on variation in product attributes across products and markets as well as their distributional characteristics (e.g., right-skewness and symmetry). If the assumption is incorrect, these instruments fail to properly capture markup variation and, as a result, have weaker identifying power. That is, when firms collude, the attributes of colluding firms’ products should be included in $\mathbf{z}^{own}$ rather than in $\mathbf{z}^{other}$.
As an illustration, consider an automobile market with two firms, Volkswagen and Toyota, and three SUVs: Atlas (A), Tiguan (T), and Highlander (H). Given that the first two SUVs are owned by Volkswagen and the third by Toyota, the ownership matrix under Bertrand-Nash competition, $\mathscr{H}^{B}$, and the ownership matrix under full collusion between the two firms, $\mathscr{H}^{C}$, are represented as follows:
$\mathscr{H}^{B}=\left(\begin{array}[]{ccc}1&1&0\\ 1&1&0\\ 0&0&1\end{array}\right)\textrm{~{}and~{}~{}}\mathscr{H}^{C}=\left(\begin{array}[]{cccc}1&1&1\\ 1&1&1\\ 1&1&1\end{array}\right).$
As derived in Appendix LABEL:app:_simple_logit, assuming that $\alpha=\beta=1$ and $X_{A}=X_{T}=X_{H}=0.2$, [^7] The market share of the outside option is 62.5%. Also, results are qualitatively robust to different values of $\alpha$, $\beta$, and $X$s. we obtain
$\left(\begin{array}[]{ccc}\frac{\partial p_{A}}{\partial x_{A}}&\frac{\partial p_{A}}{\partial x_{T}}&\frac{\partial p_{A}}{\partial x_{H}}\\ \end{array}\right)=\left\{\begin{array}[]{ccrl}(0.124&0.124&-0.037)&\textrm{if~{}}\phi=0\vspace{0.1in}\\ (0.125&0.125&0.125)&\textrm{if ~{}}\phi=1\\ \end{array}\right.$
Two points are noteworthy. First, under Bertrand-Nash competition, the price of the Atlas is more significantly affected by a change in the attributes of the Tiguan than by a change in those of the Highlander, and the two price effects are opposite in direction (0.124 vs -0.037). For instance, after improving the fuel efficiency of the Tiguan, Volkswagen would increase the prices of both the Atlas and the Tiguan to capitalize on its strengthened competitive advantage. However, the firm would lower the price of the Atlas in response to an improvement in the fuel efficiency of Toyota’s Highlander. Secondly, under full collusion, the price effect of an improvement in the Highlander’s fuel efficiency is $0.125$, which is more than three times larger in magnitude than its effect under price competition (-0.037) and aligns in direction with the effect on the Atlas’s price due to a change in the fuel efficiency of the Tiguan. In fact, the three price effects under collusion are identical in magnitude because the three SUVs are equally substitutable for one another in this example.
These observations indicate that, under price competition, the sum of own-firm product attributes (competition IVs) has greater identifying power than the sum of both own- and other-firm product attributes (collusion IVs). In contrast, if the two firms collude, collusion IVs exhibit a stronger correlation with price than competition IVs. Therefore, comparing the strength of competition IVs and collusion IVs can provide insights into whether firms in a market are competitive or collusive.
Nested logit case
To assess whether the findings above hold in a more general setting, we assume that the idiosyncratic taste shock $\varepsilon_{ijt}$ in the indirect utility function (1) follows a nested logit structure, parameterized by the nesting parameter $\sigma\in[0,1)$. [^8] Specifically, $\varepsilon_{ijt}=\zeta_{igt}+(1-\sigma)\overline{\varepsilon}_{ijt}$, where $\overline{\varepsilon}_{ijt}$ follows an i.i.d. extreme value distribution and $\zeta_{igt}$ has a unique distribution such that $\varepsilon_{ijt}$ remains extreme value distributed (Cardell97). A higher value of $\sigma$ implies stronger substitution between products belonging to the same group, indexed by $g=0,1,\ldots,G$. The outside option ($j=0$) is the sole member of group $g=0$. We then simulate equilibrium outcomes for five markets, each consisting of six firms ($F=6$) and 60 inside products ($J=60$), categorized into three groups ($G=3$). In each market, the first three firms collude ($f_{c}=1$ for $f=1,2,3$) with varying degrees of profit internalization, $\phi\in\{0,0.2,0.5,0.8,1\}$; they compete in the first market and fully collude in the last market.
Our objective is to analyze the $J\times J$ Jacobian matrix, denoted by $\frac{\partial\mathbf{p}}{\partial\mathbf{x}}$, where the $(j,k)$ entry corresponds to $\frac{\partial p_{j}}{\partial x_{k}}$ evaluated at equilibrium. Specifically, we first examine whether, under competition ($\phi=0$), the following condition holds:
$\left[\frac{\partial\mathbf{p}}{\partial\mathbf{x}}\right]_{jk}\cdot\left[\frac{\partial\mathbf{p}}{\partial\mathbf{x}}\right]_{jk^{\prime}}<0\quad\text{for $k\in\mathscr{J}_{f(j)}$ and $k^{\prime}\notin\mathscr{J}_{f(j)}$,}$
where $f(j)$ indexes the firm offering product $j$. Next, we investigate how varying the degree of profit internalization, $\phi$, affects this relationship. Analytical and computational details for deriving the Jacobian matrix are provided in Appendix LABEL:app:_nested_logit.
Figure 1: Effects of own-firm and other-firm attributes on equilibrium prices
(a) Competition ($\phi=0$)
(b) Partial profit internalization ($\phi=0.2$)
(c) Partial profit internalization ($\phi=0.5$)
(d) Partial profit internalization ($\phi=0.8$)
(e) Full profit internalization ($\phi=1$)
Notes: The figure presents heatmaps of the $J\times J$ Jacobian matrices, $\frac{\partial\mathbf{p}}{\partial\mathbf{x}}$, derived from the nested logit specification under various values of the profit internalization parameter, $\phi$. See the main text for interpretation and Appendix LABEL:app:_nested_logit for details on the data-generating process used to compute equilibrium outcomes and Jacobian matrices.
Figure 1(e) presents heatmaps of the Jacobian matrix for the five markets. Each grid represents the $(j,k)$ entry, where darker red indicates more positive values and darker blue indicates more negative values. Diagonal entries remain uncolored, as $\frac{\partial p_{j}}{\partial x_{j}}$ is significantly larger than $\frac{\partial p_{j}}{\partial x_{k}}$ for $j\neq k$. Products are indexed first by firm and then by group assignment. For instance, the first firm ($f=1$) produces nine products: the first two belong to the first group ($g=1$), the next two to the second group ($g=2$), and the remaining five to the third group ($g=3$). In each heatmap, rows and columns are marked to indicate transitions in ownership or product grouping when product $j$ differs from product $j-1$ in firm or group assignment.
The patterns of price effects in the figure align with the economic intuition derived from the simple logit case. First, in all markets, while a firm raises the prices of all its products after improving the attributes of one of its products, the direction of the price effect of other firms’ product attributes depends on industry conduct. [^9] Additionally, products with larger market shares exert stronger effects, as indicated by darker red or blue colors in the figure. That is, under competition ($\phi=0$, panel (a)), prices always decrease in response to improvements in other firms’ product attributes. In contrast, under full profit internalization ($\phi=1$, panel (e)), improvements in the attributes of products owned by colluding partners lead a firm to raise its prices with the magnitude of the price increase being larger for products in the same group.
Second, the degree of profit internalization affects how closely the impact of colluding partners’ product attributes mirrors that of a firm’s own product attributes. When profit internalization is weak ($\phi=0.2$, panel (b)), some product attributes of colluding partners have effects opposite in direction to those of the firm’s own product attributes. Specifically, a firm reduces the prices of its products that are in the same group as (and hence closer substitutes for) those of its colluding partners whose attributes are improved. [^10] Under a random coefficient utility specification where $\nu_{ir}$ in the indirect utility function (1) is heterogeneous, products with similar values of $x_{jt}^{(r)}$ serve as closer substitutes. In this case, when a firm weakly internalizes the profits of its colluding partners, it is more likely to lower the prices of products that share similar characteristics with those of its colluding partners when their attributes improve. For instance, improvements in the attributes of products in the first group ($g=1$), offered by firms 1 and 3, lead to a price reduction for firm 2’s products within the same group, as shown in panel (b), where the corresponding grids are marked in blue. [^11] Figure 1(e) also shows that the price effects are identical for products owned by the same firm and belonging to the same group, as illustrated by the long vertical strips in each heatmap. This uniformity stems from the uniform markup property under the nested logit model specification, as shown in (LABEL:eqn:_foc_nested_logit) in Appendix LABEL:app:_nested_logit; equilibrium markups are identical for products within the same group and offered by the same firm. Accordingly, in the logit case, the price effects are identical for all products owned by the same firm, as shown in Figure LABEL:fig:_heatmap_logit in the Appendix. As $\phi$ increases, the effects of colluding partners’ product attributes on prices of a firm converge to those of the firm’s own product attributes, both in direction and magnitude.
Our key findings – (i) the price effects of own-firm and other-firm attributes are opposite in direction, and (ii) the price effects of other-firm attributes reverse in direction after collusion – are consistently observed across various market structures with different values of $F$, $J$, and $G$, as well as different assignments of products to firms and groups, and varying utility and marginal cost parameters. Moreover, since these findings do not depend on imposing specific choice models, they are expected to remain valid when more flexible models are considered. Building on the analyses discussed thus far, we formalize a testing framework for detecting collusive behavior in the next section.
4 Testing framework
4.1 Proposed testing procedure
As demonstrated in Section 3 and Appendix LABEL:app:_partial_derivative, properly accounting for strong collusive firm behavior enhances the power of BLP-style instruments in the first-stage price regression. More specifically, when firms collude, incorporating the attributes of colluding partners into own-firm instruments strengthens the instruments. Conversely, if firms actually engage in price competition, this incorporation would weaken the instruments’ strength in the first-stage price regression.
Based on this intuition, we propose assessing collusive behaviors among firms by comparing the performance of IVs in two first-stage price regressions. Specifically, we construct two distinct sets of BLP-style instruments: one based on the observed firm index (competition IVs), denoted by $\mathbf{z}^{comp}$, and the other based on the suspected colluding firm index (collusion IVs), denoted by $\mathbf{z}^{coll}$. We then estimate two separate first-stage price regressions using these two sets of instruments individually and compare their relative strengths using the non-nested model selection method proposed by rivers2002model. [^12] The implementation of rivers2002model’s test in cross-sectional and panel regression settings can be inferred from wooldridge2010 (Section 13.11.2).
Our testing procedure is as follows. We first define the two BLP-style instruments for product $j\in\mathscr{J}_{ft}$ in market $t$, $\mathbf{z}_{jt}^{comp}$ and $\mathbf{z}_{jt}^{coll}$. These instruments are specified as:
$\mathbf{z}_{jt}^{comp}=\left(\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}x^{(1)}_{kt},\ldots,\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}x^{(R)}_{kt},~{}\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}\Big{(}x^{(1)}_{kt}\Big{)}^{2},\ldots,\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}\Big{(}x^{(R)}_{kt}\Big{)}^{2}\right)$ (9)
$\mathbf{z}_{jt}^{coll}=\left(\sum_{k\in\mathscr{J}_{f_{c}t}\setminus\{j\}}x^{(1)}_{kt},\ldots,\sum_{k\in\mathscr{J}_{f_{c}t}\setminus\{j\}}x^{(R)}_{kt},~{}\sum_{k\in\mathscr{J}_{f_{c}t}\setminus\{j\}}\Big{(}x^{(1)}_{kt}\Big{)}^{2},\ldots,\sum_{k\in\mathscr{J}_{f_{c}t}\setminus\{j\}}\Big{(}x^{(R)}_{kt}\Big{)}^{2}\right).$
Note that $\mathscr{J}_{f_{c}t}$ denotes the set of products owned by firm $f$ and its suspected colluding partners, as $f_{c}$ indexes these firms as a single entity. The choice of instruments and functional form in (9) is selected for brevity and clarity, with further discussion provided in Section 6.3 and Appendix LABEL:app:_functional_forms.
We emphasize that $\mathbf{z}_{jt}^{comp}$ is constructed based on the observed competitive market ownership, whereas $\mathbf{z}_{jt}^{coll}$ is constructed based on the suspected collusive market ownership. For example, suppose that the regulatory authority suspects that, among the four firms in a product market, indexed as $A$, $B$, $C$, and $D$, two firms, $A$ and $B$, are engaging in price collusion. Based on observed price and product attribute data, the authority can construct two sets of BLP-type instruments as described above. In this case, $\mathbf{z}_{jt}^{comp}$ treats firms $A$ and $B$ as two separate entities, whereas $\mathbf{z}_{jt}^{coll}$ treats these two firms as a single entity, reflecting their suspected collusion.
Next, we run the following two first-stage price regressions using $\mathbf{z}_{jt}^{comp}$ and $\mathbf{z}_{jt}^{coll}$ separately:
$p_{jt}=\gamma^{1}\cdot x_{jt}+\bm{\theta}^{1}\mathbf{z}_{jt}^{comp}+e^{1}_{jt}$ (10)
$p_{jt}=\gamma^{2}\cdot x_{jt}+\bm{\theta}^{2}\mathbf{z}_{jt}^{coll}+e^{2}_{jt},$ (11)
where $x_{jt}$ represents the exogenous characteristics of product $j$, including an intercept. Continuing with the previous example, if firms $A$ and $B$ indeed collude and internalize the profit of their colluding partner, then $\mathbf{z}_{jt}^{coll}$ should explain the equilibrium prices better than $\mathbf{z}^{comp}_{jt}$ does. On the other hand, if the two firms do not engage in collusion and all four firms compete à la Bertrand, then $\mathbf{z}_{jt}^{comp}$ should explain the equilibrium prices better.
Finally, we formalize our testing procedure by constructing the Rivers and Vuong (RV) model selection test statistic. Our RV test statistic, $T_{IV}^{RV}$, is given by:
[…]
where $\hat{Q}_{comp}$ and $\hat{Q}_{coll}$ are the averages of the sum of squared residuals of the linear regressions in (10) and (11), respectively, and $\hat{\sigma}_{IV}$ is an estimator for the asymptotic variance of $\sqrt{n}(\hat{Q}_{comp}-\hat{Q}_{coll})$. In practice, we estimate the two linear regression models in (10) and (11) and obtain the two residuals, $\hat{e}^{1}_{jt}$ and $\hat{e}^{2}_{jt}$. We then regress the difference of the two squared-residuals, $(\hat{e}_{jt}^{1})^{2}-(\hat{e}_{jt}^{2})^{2}$, on a constant term only. The $t$-statistic for this constant term corresponds to the RV test statistic given above. Clustering is easily accommodated by specifying an appropriate cluster structure in the regression of the squared-residual difference on the constant term.
It is important to note that our test allows for potential misspecification of competing models. More specifically, it does not require either model (10) or (11) to be the true pricing equation, nor do they need to represent structural pricing functions. [^13] In general, a structural pricing function cannot be estimated, as it emerges from the equilibrium behavior of firms and thus depends on numerous observed and unobserved factors. Instead, the RV test evaluates the relative fit of two alternative models, making our test both easy to implement and highly practical. This flexibility is a key advantage over other existing methods, which may require consistent estimates of the conditional mean function or structural components.
Our RV test statistic is asymptotically normal under the null hypothesis that the two models, (10) and (11), have the same fit: $Q_{comp}=Q_{coll}$. We may define two alternative hypotheses: (i) $H_{1}:Q_{comp}<Q_{coll}$ and (ii) $H_{2}:Q_{comp}>Q_{coll}$. In case of collusion, including product attributes of colluding partners in own-firm instruments is expected to enhance the strength of the instruments and improve the model fit in first-stage price regression. Therefore, a statistically significant positive test statistic (e.g., 1.65 at 0.05 significance level for a one-tailed test) can be interpreted as evidence of collusive behavior among suspected firms. In contrast, a statistically significant negative test statistic (e.g., -1.65 at the same significance level) suggests that firms engage in price competition.
4.2 Existing Testing Procedure
As elaborated in Section 2, there is a growing body of literature on firm conduct testing procedures that employ moment-based RV tests. These methods construct moments using the equilibrium markup conditions under the two alternative firm conduct models, and then compare the GMM objective functions of the two models to determine which one provides a better fit. This approach assumes that the researcher has access to a known demand system and a specified marginal cost function. Consequently, estimating demand parameters and specifying the functional form of marginal costs are necessary prerequisites. In this section, we briefly introduce the RV methodology in the existing literature. For expository purposes, we present a simplified form of this methodology, while extended discussions on this framework are provided in the footnotes accompanying the main text.
The marginal cost function is given by:
$mc_{jt}=h(x_{jt},w_{jt})+\omega_{jt},$ (13)
where $x_{jt}$ and $w_{jt}$ represent observed product attributes and cost shifters excluded from the demand function, respectively, while $\omega_{jt}$ denotes an unobserved cost component. To simplify the discussion while conveying the main idea, we impose the following two assumptions: (i) the marginal cost function, $h(x_{jt},w_{jt})$, which must be pre-specified by researchers to implement the testing procedure, is linear in $x_{jt}$ and $w_{jt}$; and (ii) marginal cost is constant with respect to the quantity produced, implying no economies of scale or scope.141414backus2021common proposed a testing procedure that incorporates a flexible specification of the cost function. Additionally, the marginal cost can be nonconstant with respect to quantities produced. In such cases, the instruments are required to satisfy additional conditions to distinguish between alternative firm conduct models. See duarte2025conduct for details and applications.
Let $D(\mathbf{x}_{t})$ denote a known demand system, where $\mathbf{x}_{t}$ represents a full set of exogenous product characteristics. Given demand estimates, the equilibrium markup for product $j$ under firm conduct model $m$ (e.g., competition, collusion, partial collusion, common ownership, etc.), $\eta_{jt}^{m}$, is also known from the profit-maximizing conditions. The marginal costs of products can then be recovered as follows:
$mc_{jt}=p_{jt}-\eta_{j}^{m}(\mathbf{S}_{t},\mathbf{p}_{t},D(\mathbf{x}_{t})).$ (14)
Under the true conduct model $m=0$, assuming no specification error, the following moment condition holds:
$E[\omega_{jt}^{0}\mathbf{z}_{jt}]=0,$
where $\omega_{jt}^{0}=p_{jt}-\eta_{jt}^{0}-h(x_{jt},w_{jt})$. A vector of excluded instruments, denoted by $\mathbf{z}_{jt}$, typically consists of BLP-style instruments (functions of exogenous product attributes) and cost shifters excluded from the demand specification, $w_{jt}$.151515backus2021common proposed a functional form for BLP-style instruments to enhance the power of the RV testing framework, drawing on the literature on optimal instruments in nonlinear GMM settings (chamberlain1987asymptotic). This functional form results in a scalar moment that is independent of the choice of the weighting matrix used to form the objective function. Additionally, the choice of instruments (beyond BLP-style instruments) to satisfy falsifiable conditions to distinguish between two alternative conduct models affects the validity of testing procedure. See dearing2024learning for discussions on falsifiable conditions and empirical applications.
The existing methods choose between two non-nested conduct models, $m=1$ and $m=2$. More specifically, the linear parameters in the marginal cost function $h(x_{jt},w_{jt})$ are estimated by running an OLS regression of the marginal cost under conduct model $m$:
$p_{jt}-\eta_{jt}^{m}=h(x_{jt},w_{jt})+\omega_{jt}^{m},$
[…]
$\hat{Q}^{m}(\eta^{m})=\hat{g}_{m}^{\prime}\hat{W}\hat{g}_{m},$
where, given the total number of observations $n$,
$\hat{g}_{m}=n^{-1}\sum_{t}\sum_{j}\mathbf{z}_{jt}^{\prime}\widehat{\omega_{jt}^{m}}\quad\textrm{and}\quad\hat{W}=n\cdot\left(\sum_{t}\sum_{j}\mathbf{z}_{jt}\mathbf{z}_{jt}^{\prime}\right)^{-1}.$
Finally, the RV test statistic based on equilibrium markup conditions, $T_{markup}^{RV}$, is given by:
[…]
where $\hat{\sigma}_{markup}$ is an estimator for the asymptotic variance of $\sqrt{n}(\hat{Q}_{1}(\eta^{1})-\hat{Q}_{2}(\eta^{2}))$. The exact form of $\hat{\sigma}_{markup}$, along with a detailed discussion on adjustments for clustering and two-step estimation errors arising from demand estimation, is provided in duarte2024testing. This RV statistic can be computed using the Python package PyRVtest (pyrvtest). [^17] Additionally, diagnostics for weak instruments in this testing framework, as illustrated in (duarte2024testing), can be executed using this package.
4.3 Summary
The basic idea behind our proposed conduct testing procedure aligns with that of existing methods: to select the firm conduct model that best fits the observed market outcomes. In our approach, model fit is evaluated using the first-stage price regression, whereas existing approaches rely on the GMM objective function. Unlike these methods, our testing procedure does not require demand estimation or marginal cost specification, making it straightforward to implement. Furthermore, our procedure is data-efficient, as it can be performed without market share data or additional instruments (beyond BLP-style instruments), including exogenous cost shifters. As demonstrated in the following section, the performance of our testing method is comparable to that of existing approaches. Consequently, our method complements these approaches by serving as a preliminary diagnostic tool for assessing industry conduct.
5 Monte Carlo set-up
5.1 Data-generating process
[…]
The indirect utility of consumer $i$ in market $t$ from consuming product $j$, and the marginal cost of product $j$ in market $t$, are specified as follows:
$u_{ijt}=\beta_{1}+\beta_{2}x_{jt}-\alpha p_{jt}+\xi_{jt}+\sigma_{x}v_{i}x_{jt}+\varepsilon_{ijt},$ (16)
$mc_{jt}=\gamma_{1}+\gamma_{2}x_{jt}+\gamma_{3}w_{jt}+\omega_{jt}.$ (17)
There are two exogenous product characteristics – a constant term and $x_{jt}$ – and one exogenous cost shifter, $w_{jt}$, where both $x_{jt}$ and $w_{jt}$ are randomly drawn from a standard uniform distribution. Price, $p_{jt}$, is an endogenous product characteristic determined by equilibrium conditions in a differentiated goods market under a Bertrand-Nash framework. The true linear demand parameters are given by $\alpha=1$, $\beta_{1}=-4.5$, and $\beta_{2}=6$. When included, the heterogeneous component of demand is given by $\sigma_{x}v_{i}x_{jt}$, where $\sigma_{x}=3$ and $v_{i}$ follows a standard normal distribution. Cost parameters are given by $\gamma_{1}=2$, $\gamma_{2}=1$, and $\gamma_{3}=0.2$. The unobservable components of demand and cost, $(\xi_{jt},\omega_{jt})$, are randomly drawn from a mean-zero bivariate normal distribution with standard deviations $\sigma_{\xi}=0.2$, $\sigma_{\omega}=0.2$, and covariance $\sigma_{\xi\omega}=0.1$.
Firms may engage in collusion, internalizing the profits of their colluding partners as part of their own. The degree of collusion is measured by $F/F_{c}$, where $F_{c}\in\{1,\ldots,F-2,F-1\}$ denotes the effective number of competitors – i.e., the number of distinct competitive entities after treating all colluding firms as a single entity. We assume that the first $F-F_{c}+1$ firms either collude or are suspected of colluding by the regulatory authority. For example, suppose $F=6$ and $F_{c}=4$. In this case, the degree of collusion is 1.5, with the first three firms ($f=1,2,3$) colluding or suspected of colluding to jointly set profit-maximizing prices, leaving the effective competitors indexed as $f_{c}=1,4,5,6$. The degree of profit internalization, represented by the conduct parameter $\phi\in[0,1]$, quantifies the weight assigned to the profits of colluding partners, as shown in equation (4). Continuing with the previous example, when $F_{c}=4$ but $\phi=0$, this configuration represents a scenario in which regulatory authorities falsely suspect three firms of collusion when, in fact, they are engaged in price competition.
Equilibrium prices in each market are endogenously determined by the Bertrand first-order conditions (5). These prices are solved using the fixed-point algorithm proposed by morrow2011fixed and implemented in PyBLP (conlon2020best). We use nine Gauss-Hermite quadrature nodes to numerically integrate the choice probabilities when a random coefficient is included in the indirect utility function.
We also consider the case in which the product attribute is correlated with the unobservable demand component, $\xi_{jt}$, making it endogenous. We denote the endogenous attribute as $x_{jt}^{endo}$ and derive it as follows. First, we compute $x_{jt}^{unscaled}=x_{jt}+\rho\cdot\xi_{jt}$, where $\rho\in\{-10,-5,-1,0,1,5,10\}$ determines the degree and direction of endogeneity; the case $\rho=0$ corresponds to an exogenous product attribute. Next, given that the variation of $x_{jt}^{unscaled}$ depends on $\rho$, we apply min-max normalization to define $x_{jt}^{endo}=\frac{x_{jt}^{unscaled}-x_{t}^{\min}}{x_{t}^{\max}-x_{t}^{\min}}$, where $x_{t}^{\min}$ and $x_{t}^{\max}$ are the minimum and maximum values of $x_{jt}^{unscaled}$ in market $t$. In this way, we fix the support of the endogenous attribute: $x_{jt}^{endo}\in[0,1]$ for any $\rho$.
Each Monte Carlo configuration is represented by a unique five-tuple $(J,F,T,\phi,F_{c})$. Specifically, $J=36$, $F\in\{1,2,\ldots,36\}$, $T\in\{10,100\}$, $\phi\in\{0,0.1,\ldots,1\}$, and $F_{c}\in\{1,2,\ldots,F-1\}$. [^19] When the DGP includes an endogenous product attribute, $\phi$ is fixed at 1, and $\rho$ is varied to generate unique five-tuples $(J,F,T,\rho,F_{c})$. For each configuration, we generate 500 simulated datasets ($S=500$) to evaluate the finite-sample performance of the proposed test statistic, $T_{IV}^{RV}$, and compare it with the performance of the test based on equilibrium markup conditions, $T_{markup}^{RV}$. We also investigate the performance of $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as instruments for demand estimation. Furthermore, in Section 7, we focus on cases with varying $F$ and fixed $J$ but without collusive behavior ($\phi=0$ and $F_{c}=F$) to evaluate the performance of BLP-type IVs, following the approach of armstrong2016large.
5.2 Computation of test statistics
For each dataset with configuration $(J,F,T,\phi,F_{c})$, we construct $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as outlined in Section 4.1.202020Note that the summation of the constant term, often referred to as the product counts IV, cannot be utilized in our DGP setting because it fails to generate cross-firm or cross-market variations in the number of products. Moreover, the inclusion of second-order polynomial IVs in (9) ensures that the rank condition for the GMM framework is satisfied in our DGP when a random coefficient is incorporated into the indirect utility function (16). Note that even when $\phi=0$, we still construct $\mathbf{z}^{coll}$ based on a scenario in which the researcher falsely suspects collusion among the firms of interest. To assess the relevance of these two instruments under a given true conduct model, we compute two F-statistics associated with (i) $\bm{\theta}^{1}=0$ (Fstat1) and (ii) $\bm{\theta}^{2}=0$ (Fstat2) from equations (10) and (11), respectively. Our proposed test statistic, $T_{IV}^{RV}$, is then calculated using equation (12) and the description provided therein. Heteroskedasticity-robust standard errors are used to compute $\hat{\sigma}_{IV}$ in equation (12).
To compute $T_{markup}^{RV}$, we first estimate the demand parameters using the standard own- and other-firm IVs, defined as:
$\mathbf{z}_{jt}^{own}=\left(\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}x_{kt},\sum_{k\in\mathscr{J}_{ft}\setminus\{j\}}x_{kt}^{2}\right),\quad\mathbf{z}_{jt}^{other}=\left(\sum_{k\in\mathscr{J}_{t}\setminus\mathscr{J}_{ft}}x_{kt},\sum_{k\in\mathscr{J}_{t}\setminus\mathscr{J}_{ft}}x_{kt}^{2}\right),$ (18)
as excluded instruments $\mathbf{z}_{jt}=\left(\mathbf{z}_{jt}^{own},\mathbf{z}_{jt}^{other}\right)$. Note that these instruments are constructed based on observed ownership rather than suspected ownership. Next, we specify the marginal cost function as a linear function of the constant term and $x_{jt}$ only, even though it is also linear in a cost shifter $w_{jt}$, as shown in equation (17). [^21] Accordingly, we do not include $w_{jt}$ as an instrument when estimating the demand. Thus, we consider a scenario in which the cost function is potentially misspecified, either due to data limitations on cost shifters or the omission of a relevant cost shifter. This approach enables us to evaluate the performance of the two testing procedures without using additional data. We then follow the procedure outlined in Section 4.2 to construct $T_{markup}^{RV}$ under two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full profit internalization under industry conduct consistent with the effective firm index used in the DGP). [^22] Although the testing procedure built on equilibrium markup conditions allows us to compare conduct models with other values of $\phi$, such as $\phi=0.2$ versus $\phi=0.8$, the primary interest of researchers and regulatory authorities is often to test full profit internalization ($\phi=1$) against competition ($\phi=0$). Therefore, we compute $T_{markup}^{RV}$ under these two conduct models. The standard error $\hat{\sigma}_{markup}$ in equation (15) is adjusted for two-step demand estimation errors and is heteroskedasticity-robust. The computation of $T_{markup}^{RV}$ is done using PyRVtest (pyrvtest).
6 Monte Carlo evidence: testing performance of $T_{IV}^{RV}$ vs $T_{markup}^{RV}$
6.1 Power of instruments and estimation performance
We begin by assessing the power of our proposed testing procedure outlined in Section 4.1. To do so, we consider a scenario with four firms, two of which collude ($F=4,F_{c}=3$). We generate Monte Carlo datasets, following the design in Section 5, for various values of $\phi\in[0,1]$. The top panel of Table 1 presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel shows the results when it is included.
[…]
Table 1: Comparison of IV performance: $\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$ ($F=4$, $F_{c}=3$, $T=100$)
$\mathbf{z}^{comp}$ $\mathbf{z}^{coll}$
Median values ($S=500$) of % of Fstat2 Median
$\phi$ $|\alpha-\hat{\alpha}|$ $|\sigma_{x}-\hat{\sigma}_{x}|$ $\text{Fstat}_{1}$ $|\alpha-\hat{\alpha}|$ $|\sigma_{x}-\hat{\sigma}_{x}|$ $\text{Fstat}_{2}$ $>$ Fstat1 of $s_{o}$
Panel A: without a random coefficient
0 (competition) 0.143 10.377 0.190 6.123 0.060 0.659
0.1 0.143 10.329 0.171 7.215 0.132 0.660
0.2 0.142 10.308 0.155 8.663 0.266 0.662
0.3 0.143 10.296 0.144 10.377 0.472 0.663
0.4 0.144 10.269 0.134 12.557 0.694 0.664
0.5 0.144 10.209 0.121 15.164 0.866 0.665
0.6 0.144 10.155 0.108 18.336 0.958 0.666
0.7 0.146 10.213 0.097 21.690 0.990 0.667
0.8 0.146 10.177 0.090 25.504 1.000 0.668
0.9 0.146 10.199 0.083 29.670 1.000 0.669
1 0.148 10.095 0.077 34.193 1.000 0.669
Panel B: with a random coefficient
0 (competition) 0.552 1.863 25.102 0.727 1.148 10.861 0.000 0.604
0.1 0.549 1.867 24.698 0.545 0.920 13.181 0.002 0.606
0.2 0.541 1.770 24.414 0.321 0.620 16.456 0.056 0.607
0.3 0.536 1.698 23.962 0.221 0.474 20.728 0.228 0.609
0.4 0.536 1.707 23.853 0.167 0.397 26.078 0.590 0.610
0.5 0.537 1.711 23.642 0.132 0.367 32.597 0.894 0.612
0.6 0.533 1.613 23.038 0.109 0.325 40.129 0.986 0.613
0.7 0.520 1.552 22.737 0.091 0.312 48.787 1.000 0.615
0.8 0.504 1.472 22.420 0.079 0.310 58.424 1.000 0.616
0.9 0.510 1.495 21.959 0.070 0.300 69.041 1.000 0.618
1 0.509 1.434 21.358 0.063 0.291 80.495 1.000 0.619
Notes: The table compares the median absolute errors of the estimated price and nonlinear coefficients, as well as the median F-statistics, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=4,T=100,\phi,F_{c}=3$), obtained using $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as instruments individually. The top panel presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel presents the results when it is included.
Our simulation results presented in Table 1 also show that the median absolute error of the estimated price coefficient $\hat{\alpha}$ decreases significantly in response to a rise in $\phi$ when using $\mathbf{z}^{coll}$ as instruments. As a result, $\mathbf{z}^{coll}$ outperforms $\mathbf{z}^{comp}$ in estimating $\alpha$ when firms collude, even with a low degree of profit internalization ($\phi\geq 0.4$ without a random coefficient and $\phi\geq 0.1$ with a random coefficient).
A similar pattern is observed for the median absolute error of the estimated non-linear coefficient $\hat{\sigma}_{x}$. Moreover, for any $\phi$, the median absolute error obtained using $\mathbf{z}^{coll}$ as instruments is smaller than that obtained using $\mathbf{z}^{comp}$. This finding is not contradictory to our earlier observation that, when the degree of profit internalization is low ($\phi\leq 0.3$), using $\mathbf{z}^{coll}$ instead of $\mathbf{z}^{comp}$ leads to lower first-stage F-statistics. This is because identifying $\sigma_{x}$ requires instruments that are correlated not only with price but also with market shares. [^23] As a simple illustration, consider a nested logit specification, a type of random coefficient model. After inverting a demand system (Berry94), the resulting estimation equation using a standard linear instrumental variables approach is as follows: $\log\frac{s_{jt}}{s_{ot}}=\beta x_{jt}-\alpha p_{jt}+\sigma\log s_{j|g,t}+\xi_{jt},$ where $s_{j|g,t}$ denotes the market share of product $j$ within group $g$. Identifying this random coefficient model requires instruments for both the endogenous price and the endogenous market share $s_{j|g,t}$. See berry2014identification for the instrument requirements under general nonparametric settings. We revisit this point in Section 7, where we discuss the role of own- and other-firm instruments.
In sum, taking the true industry conduct into consideration when constructing instruments increases the strength of the instruments and enhances the estimation performance of the demand parameters.
6.2 Comparison of Testing Power
Now, we compute and compare the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$. Specifically, we investigate how the degree of profit internalization, $\phi$, as well as the degree of collusion, $F/F_{c}$, influence these two statistics. Recall that for both statistics, a statistically significant positive value (greater than 1.65 at the 0.05 significance level) indicates collusion, while a statistically significant negative value (less than -1.65 at the same significance level) indicates competition.
Table 2 reports these two statistics under various Monte Carlo configurations elaborated in Section 5.1. We first examine the results without a random coefficient presented in the upper panel. Overall, our test statistic, $T_{IV}^{RV}$, performs relatively well compared to the existing statistic, $T_{markup}^{RV}$, in detecting collusion. In contrast, the latter tends to outperform the former in rejecting collusion in favor of competition at lower values of $\phi$. For example, when all six firms collude ($F_{c}=1$) and fully internalize the profits of other firms ($\phi=1$), both tests reject the competition hypothesis: $T_{IV}^{RV}=3.309$ and $T_{markup}^{RV}=3.564$. As the degree of profit internalization decreases, $T_{IV}^{RV}$ also declines but remains statistically significant (and positive) for $\phi\geq 0.7$, while $T_{markup}^{RV}$ remains statistically significant only for $\phi\geq 0.9$. For $\phi=0.1$, only $T_{markup}^{RV}$ has a statistically significant negative value (-6.088), allowing the rejection of collusion. When $F_{c}=1$, there is no within-market variation in the collusion IVs, and their identification power relies solely on cross-market variations. While this suggests that the performance of our test may improve as the number of available markets increases, our testing procedure performs even better when $F_{c}\in\{2,3,4\}$ compared to the $F_{c}=1$ scenario in detecting collusive behaviors among firms across a broader range of $\phi$.
Table 2: (Exogenous) $F=6$ and $T=100$
$T_{IV}^{RV}$ ($\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$) $T_{markup}^{RV}$ ($\phi=0$ vs $\phi=1$)
$\phi$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$
Panel A: without a random coefficient
0 (competition) -1.657 -1.547 -1.483 -1.299 -0.925 -7.133 -3.872 -2.143 -1.033 -0.385
0.1 -1.458 -1.012 -1.030 -0.978 -0.758 -6.088 -3.360 -1.839 -0.880 -0.326
0.2 -1.112 -0.063 -0.129 -0.374 -0.451 -5.106 -2.819 -1.524 -0.719 -0.271
0.3 -0.604 0.992 0.845 0.328 -0.144 -4.083 -2.228 -1.175 -0.556 -0.227
0.4 0.011 1.986 1.782 1.004 0.211 -2.930 -1.614 -0.820 -0.393 -0.165
0.5 0.650 2.911 2.620 1.672 0.574 -1.758 -0.958 -0.453 -0.221 -0.131
0.6 1.272 3.787 3.405 2.294 0.886 -0.644 -0.300 -0.081 -0.055 -0.075
0.7 1.849 4.607 4.167 2.875 1.214 0.477 0.380 0.277 0.119 -0.011
0.8 2.372 5.373 4.880 3.404 1.522 1.553 1.010 0.593 0.305 0.043
0.9 2.851 6.111 5.599 3.925 1.825 2.597 1.622 0.944 0.469 0.097
1 3.309 6.824 6.277 4.414 2.105 3.564 2.208 1.283 0.643 0.162
Panel B: with a random coefficient
0 (competition) -2.737 -2.867 -2.905 -2.716 -2.039 -8.601 -2.536 -1.480 -0.754 -0.383
0.1 -2.636 -2.343 -2.344 -2.257 -1.683 -8.092 -2.353 -1.322 -0.640 -0.318
0.2 -2.549 -0.929 -0.963 -1.239 -1.198 -7.535 -2.126 -1.158 -0.512 -0.255
0.3 -2.460 0.845 0.676 -0.030 -0.612 -6.894 -1.844 -0.944 -0.354 -0.181
0.4 -2.320 2.644 2.302 1.168 0.019 -6.091 -1.504 -0.676 -0.200 -0.083
0.5 -2.109 4.343 3.826 2.293 0.639 -5.143 -1.090 -0.412 -0.037 0.004
0.6 -1.659 5.945 5.252 3.337 1.219 -4.004 -0.559 -0.036 0.110 0.071
0.7 -0.732 7.466 6.606 4.295 1.789 -2.480 0.009 0.276 0.280 0.149
0.8 0.721 8.896 7.892 5.220 2.307 -0.971 0.686 0.556 0.428 0.199
0.9 2.290 10.225 9.136 6.095 2.784 -0.833 1.299 0.818 0.548 0.271
1 3.620 11.442 10.300 6.911 3.227 -0.459 1.738 1.075 0.708 0.338
Notes: The table reports the median values of the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=6,T=100,\phi,F_{c}$). $T_{markup}^{RV}$ is constructed under the two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full profit internalization under industry conduct consistent with the effective firm index). The top panel presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel presents the results when it is included.
The relative testing power of our method is higher when a random coefficient is included in the consumer utility function. [^24] This higher testing power suggests that overall, the price effect of product attributes is more sensitive to industry conduct under the random coefficient specification than under the logit specification employed in our Monte Carlo study. For any degree of collusion ($F/F_{c}$), it effectively detects collusion at high values of $\phi$ while rejecting it in favor of competition at lower values of $\phi$. Conversely, the inclusion of a random coefficient tends to reduce the testing power of $T_{markup}^{RV}$. These results highlight the advantage of our approach: it performs comparably to, or outperforms, the existing method under various Bertrand-Nash collusive scenarios, while its implementation does not require estimating the demand and supply system.
Figure LABEL:fig:_p-value_iv_vuong in the Appendix presents the median p-values under the alternative hypothesis, $H_{2}:Q_{comp}>Q_{coll}$, derived from extensive additional simulations. Clearly, there is a positive monotonic relationship between testing power and the degree of profit internalization across all $(F,F_{c})$ configurations considered in our Monte Carlo study. The degree of collusion ($F/F_{c}$) also influences testing power; except in cases where all firms in the market collude, a higher degree corresponds to greater power in detecting collusive behavior among firms.
So far, we have treated the product attribute, $x_{jt}$, as an exogenous variable. However, this assumption may not hold in situations where firms determine product attributes and prices simultaneously with the contemporaneous unobservable components of demand, $\xi_{jt}$. In such cases, BLP-style instruments fail to satisfy the exclusion restriction, rendering them invalid and leading to inconsistent demand estimates. To evaluate the testing power of the two statistics under this model misspecification, we generate endogenous product attributes as described in Section 5.1. The true profit internalization parameter, $\phi$, is fixed at 1, while the direction and degree of endogeneity are parameterized by $\rho\in\{-10,-5,-1,0,1,5,10\}$ as explained in Section 5.1.
The top panel of Table 3 reports the results obtained without a random coefficient in the indirect utility specification (16) in the DGP. While the power of both test statistics remains robust across various endogeneity scenarios, the direction of correlation between $\xi_{jt}$ and $x_{jt}$ affects the testing power. When the correlation is negative, the testing power declines; conversely, when the correlation is positive, the testing power improves. Results obtained with a random coefficient in the utility specification, presented in the bottom panel of the table, indicate that the testing power of both statistics tends to be higher when the model is misspecified. [^25] We also consider an alternative DGP for the endogenous product attribute. Specifically, the triplet $(x_{jt}^{endo},\xi_{jt},\omega_{jt})$ is drawn from a mean-zero trivariate normal distribution with a covariance matrix: $\begin{bmatrix}0.2&\rho&0\\ \rho&0.2&0.1\\ 0&0.1&0.2\end{bmatrix},$ where $\rho$ is in between -0.2 and 0.2. The results remain qualitatively unaffected.
In sum, our results are robust to model misspecification, which is one of the key advantages of the RV test, as highlighted by duarte2024testing. Moreover, our test statistic, $T_{IV}^{RV}$, exhibits higher testing power than the existing statistic, $T_{markup}^{RV}$, across all specifications except in the case of $F_{c}=1$, where the difference is only marginal.
Table 3: (Endogenous) $F=6$ and $T=100$
$T_{IV}^{RV}$ ($\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$) $T_{markup}^{RV}$ ($\phi=0$ vs $\phi=1$)
$\rho$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$
Panel A: without a random coefficient
Exogenous
$\rho=0$ 3.546 7.240 6.630 4.679 2.268 3.905 2.469 1.489 0.739 0.243
Endogenous ()(-)( - )
$\rho=-1$ 2.523 5.611 5.195 3.600 1.698 2.853 1.825 1.053 0.528 0.193
$\rho=-5$ 2.446 5.644 5.231 3.650 1.663 2.782 1.755 1.011 0.484 0.139
$\rho=-10$ 2.440 5.658 5.231 3.648 1.662 2.818 1.763 1.038 0.489 0.111
Endogenous (+)(+)( + )
$\rho=1$ 4.494 7.743 7.065 4.984 2.431 4.858 3.060 1.853 0.967 0.366
$\rho=5$ 4.849 8.185 7.457 5.261 2.573 5.253 3.327 2.022 1.110 0.492
$\rho=10$ 4.857 8.193 7.488 5.271 2.606 5.130 3.195 1.913 0.979 0.359
Panel B: with a random coefficient
Exogenous
$\rho=0$ 3.650 11.492 10.361 6.948 3.248 0.265 1.712 1.033 0.720 0.326
Endogenous ()(-)( - )
$\rho=-1$ 4.005 11.889 10.812 7.294 3.407 3.953 2.047 1.212 0.697 0.321
$\rho=-5$ 4.057 12.188 11.114 7.466 3.480 4.839 2.061 1.145 0.677 0.297
$\rho=-10$ 4.065 12.180 11.129 7.494 3.490 4.907 2.090 1.173 0.709 0.308
Endogenous (+)(+)( + )
$\rho=1$ 5.126 12.991 11.749 7.835 3.709 5.299 2.426 1.319 0.894 0.421
$\rho=5$ 5.446 13.406 12.234 8.215 3.901 6.463 2.499 1.380 0.835 0.406
$\rho=10$ 5.441 13.434 12.277 8.232 3.906 6.502 2.424 1.348 0.847 0.358
Notes: The table reports the median values of the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=6,T=100,\rho,F_{c}$). The product attribute $x_{jt}$ is treated as an endogenous variable. The direction and degree of endogeneity are parameterized by $\rho\in\{-10,-5,-1,0,1,5,10\}$, while the true profit internalization parameter $\phi$ is fixed at 1. $T_{markup}^{RV}$ is constructed under the two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full profit internalization under industry conduct consistent with the effective firm index). The top panel presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel presents the results when it is included.
Our simulation setting so far has assumed a scenario in which researchers have sufficient observations from many markets ($T=100$). To evaluate the performance of the two testing procedures when only a few markets are available in the data, we repeat our analysis with $T=10$. Results summarized in Tables LABEL:tab:_exo_vuong_few and LABEL:tab:_endo_vuong_few in the Appendix reveal that, as expected, both statistics experience a substantial loss in testing power under this data limitation. Interestingly, $T_{IV}^{RV}$ continues to yield significantly positive statistics across various collusive scenarios, whereas $T_{markup}^{RV}$ retains significant power only under a few specific configurations: positive endogeneity of $x_{jt}$ ($\rho\geq 5$), $F_{c}=1$, and the inclusion of a random coefficient in the consumer utility specification, as shown in the bottom panel of Table LABEL:tab:_endo_vuong_few. The results also indicate that $T_{markup}^{RV}$ tends to reject collusion more often than $T_{IV}^{RV}$ when the true industry conduct is closer to competition (low $\phi$).
6.3 Alternative functional forms for BLP-style instruments
For expositional purposes, we have used only one functional form for BLP-style instruments: the summation of other product attributes (summation IVs), as originally proposed by Berry95. To assess the robustness of our results to alternative functional forms, we incorporate Differentiation IVs (gandhi2019measuring) up to third-order polynomials in our Monte Carlo study.
Our results largely align with previous findings. Notably, our test statistic, $T_{IV}^{RV}$, whose testing power remains relatively unaffected by changes in functional forms, continues to outperform $T_{markup}^{RV}$ in detecting collusive behavior among firms across various collusive scenarios. Interestingly, the testing power of $T_{markup}^{RV}$ declines substantially when only Local and Quadratic IVs are used but is restored when the third-order analogue is incorporated within Differentiation IVs. We also observe similar effects of model misspecification arising from the endogenous product attribute $x_{jt}$ to those presented in Table 3. Additionally, the use of Differentiation IVs enhances the accuracy of estimating the non-linear coefficient $\sigma_{x}$, consistent with the findings of gandhi2019measuring. Further details on the alternative functional form design and extensive discussions of the results are provided in Appendix LABEL:app:_functional_forms.
The testing power of $T_{IV}^{RV}$ depends on the strength of the instruments in the first-stage price regressions. We use summation IVs as the baseline instruments for our proposed testing procedure, primarily because, compared to other specifications, they provide stronger identification power in the first-stage price regression across most Monte Carlo configurations considered. [^26] Specifically, the median values of first-stage F-statistics obtained using summation IVs are higher than those from Differentiation IVs across various Monte Carlo specifications, including those presented in Section 7. The core of our approach lies not in selecting the optimal functional forms for BLP-style instruments but in leveraging alternative firm indexes based on the observed and suspected ownership structures when constructing these instruments.
6.4 Discussion
The most notable difference between our proposed testing procedure and the existing one is that ours does not require researchers to estimate the demand system or impose a marginal cost specification. Therefore, $T_{IV}^{RV}$ does not rely on the statistical properties of demand estimates, circumventing potential challenges associated with estimating the model by non-linear GMM and the inputs required for this task. [^27] Moreover, our approach is expected to be less dependent on the functional form of choice probabilities – such as nested logit, multivariate probit, and multiple-choice models – compared to existing methods that require demand estimation as a prerequisite. This is because our testing procedure relies solely on firm indexing and the nature of non-nested tests. In contrast, reliable and powerful testing with $T_{markup}^{IV}$ depends on obtaining consistent and efficient demand estimates. Furthermore, since our testing method simply compares the model fit of two linear price regressions, it can be implemented in data-limited settings where researchers lack information on product market shares. Given these properties, our method can serve as a preliminary tool for researchers and regulatory authorities to diagnose collusive behavior among suspected firms in the market.
Once preliminary results indicate the presence of collusion, researchers can proceed with existing tests on industry conduct. The advantage of the existing statistic, $T_{markup}^{RV}$, is that researchers can test any model of conduct, provided the first-order profit-maximizing conditions can be derived from the model, and determine which model best fits the observed data based on the moment restrictions imposed on the supply-side model. On the other hand, our approach, while not assuming any particular supply model, aims to detect price collusion among firms grounded in Bertrand-Nash equilibrium conditions in differentiated goods markets. Therefore, these two approaches can complement each other, creating a more practical, efficient, and powerful framework for testing firm conduct.
In addition to its preliminary role in detecting collusive behavior, our proposed testing procedure can also be used to construct stronger BLP-style instruments. For instance, once researchers observe evidence of collusive behavior, they can construct $\mathbf{z}^{coll}$ based on the colluding firm index and use it in demand estimation instead of $\mathbf{z}^{comp}$. Improved performance in the first-stage demand estimation, achieved through the use of instruments that properly capture true firm conduct behavior, would ultimately enhance performance in the second-stage estimation, where (feasible) optimal instruments are employed (e.g., reynaert2014improving, gandhi2019measuring, and conlon2020best).
7 The role of own- and other-firm IVs in demand estimation
Our proposed testing method for firm conduct utilizes $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$, which are functions of product attributes produced by the firm and its suspected colluding partners (e.g., summation of attributes or squared attributes). Here, we set aside the issue of collusion and focus on the role of the standard own- and other-firm instruments, $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$, in demand estimation. The simulation study presented in this section provides additional intuition and aids in interpreting the previous simulation results related to our proposed testing procedure.
The Monte Carlo setup is as follows: We fix the number of products at $J=36$ and vary the number of firms $F\in\{1,2,\ldots,36\}$. Since we do not consider any form of collusive behavior among firms, $\phi$ is fixed at 0. We generate instruments $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$ as defined in equation (18). To investigate their roles in demand estimation, we report the median F-statistics from the first-stage price regression and the median absolute errors of the estimated utility parameters across 500 simulations under the following instrument configurations: (i) $\mathbf{z}^{own}$ only, (ii) $\mathbf{z}^{other}$ only, and (iii) $\mathbf{z}^{both}=(\mathbf{z}^{own},\mathbf{z}^{other})$.
The results for the utility specification without (with) a random coefficient, illustrated in the upper panel (bottom panel, respectively) of Figure 2, reveal several interesting points. [^28] Tables LABEL:tab:_own_vs_other_full and LABEL:tab:_own_vs_other_full_rc in the Appendix provide the full numeric results, including the median root-mean-squared errors (RMSE) of the estimated coefficients. First, own-firm instruments exhibit greater identification power than other-firm instruments in the first-stage price regression, as shown in Figures 2(a) and (c). Consequently, the median absolute error of the estimated price coefficient is lower when $\mathbf{z}^{own}$ is used as the instrument, as shown in Figures 2(b) and (d). These observations are consistent with the argument in Section 3 and previous literature (Berry95; gandhi2021empirical) that firms internalize the cross-price elasticities of their own products when setting profit-maximizing prices, resulting in greater explanatory power of own-firm product attributes in determining equilibrium prices. They also align with our simulation results presented in the previous section and echo the importance of correctly indexing firms based on the true conduct model.
In the previous sections, we designed our Monte Carlo study and conducted simulations using only own-firm instruments, $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$, as they exhibit greater identification power than other-firm instruments in the first-stage price regression. In fact, our testing procedure can be extended to incorporate other-firm instruments constructed in the same manner as the own-firm instruments, based on the original firm and the suspected colluding firm indices. Excluding the product attributes of colluding partners from other-firm instruments enhances their identification power because, as outlined in Section 3, the price effects of colluding and non-colluding firms’ product attributes are opposite in direction. [^29] Simulation results available from the authors upon request are both quantitatively and qualitatively similar to those obtained using only the own-firm instruments. Additionally, it is worth noting that when testing whether all firms in the market collude ($F_{c}=1$), constructing other-firm instruments under the suspected colluding firm index is impossible, making it more likely for our testing procedure to diagnose that the model of competition provides a better fit.
Figure 2: Own-firm vs other-firm instruments
Panel A: without a random coefficient
(a) Median First stage F-stat
(b) Median $|\hat{\alpha}-\alpha|$
Panel B: with a random coefficient
(c) Median First stage F-stat
(d) Median $|\hat{\alpha}-\alpha|$
(e) Median $|\hat{\sigma}_{x}-\sigma_{x}|$
Notes: The figure shows the median F-statistic from the first-stage price regression and the median absolute errors of the estimated utility parameters, $\hat{\alpha}$ and $\hat{\sigma}_{x}$, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F,T=100$). The upper panel presents results for the utility specification without a random coefficient, while the bottom panel shows results with a random coefficient.
Second, the bottom panel of the figure shows that while own-firm instruments are more effective in the first-stage price regression and more powerful for identifying the price coefficient, other-firm instruments are more useful for identifying the non-linear coefficient $\sigma_{x}$. [^30] Using $\mathbf{z}^{other}$ instead of $\mathbf{z}^{own}$ consistently results in a lower median absolute error and a lower median RMSE for the estimated non-linear coefficient, as shown in Table LABEL:tab:_own_vs_other_full_rc in the Appendix. These observations help explain the results in the bottom panel of Table 1, where using $\mathbf{z}^{coll}$ leads to a smaller median absolute error for the estimated non-linear coefficient, even when the degree of profit internalization is low. Specifically, variations in product attributes of other firms unintentionally included in $\mathbf{z}^{coll}$ may assist in identifying the non-linear coefficient.
Third, the less concentrated the market, the lower the explanatory power of the two BLP-style instruments. This occurs because, as competition among firms intensifies, equilibrium markups converge, which in turn weakens the correlation between markups and instruments. For instance, assuming a simple logit demand, when the market is competitive, the values inside the bracket in equation (8) are not significantly different across products in the market. In fact, this finding aligns with the observation made by armstrong2016large in the extreme case of $J\rightarrow\infty$. [^31] Specifically, Armstrong (2016) examined cases where markets consist of many single-product firms ($J_{t}=F_{t}$). He demonstrated that as $J_{t}\to\infty$, the explanatory power of BLP-style instruments for equilibrium markups diminishes. Cross-firm or market-level variations in equilibrium markups can help maintain identification power in such cases. See Armstrong (2016) for further details. On the other hand, in a more concentrated market, greater variation in markup sizes reduces the sensitivity of the correlation between markups and instruments to fluctuations from unobserved components.
8 Empirical application
In this section, we apply our testing method to study industry conduct in two differentiated product markets in South Korea: the new passenger car market and the instant noodles market. We then examine whether the results align with anecdotal and analytical evidence of industry conduct in each market. Namely, automobile manufacturers owned by the same parent company jointly determine the prices of their car models, while instant noodle prices remain significantly below the collusive level despite suspicions from the Korean Fair Trade Commission (KFTC).
8.1 The new passenger car market
South Korea is one of the largest car markets in the world, with approximately 1.6 million vehicles sold in 2022.323232https://www.statista.com/statistics/265891/vehicles-sales-in-selected-countries/ The two largest firms, Hyundai and Kia, together accounted for nearly 70 percent of total vehicle sales over the past decade. Since both are owned by the Hyundai Motor Group, following Kia’s merger into the group in 1998, pricing is likely coordinated between the two brands. [^33] We use the terms “firm” and “brand” interchangeably. We test this joint profit maximization against own-profit maximization, in which each firm sets prices independently. We also consider three alternative firm conduct scenarios: joint profit maximization among German brands; collusion among domestic automakers and among foreign automakers; and full collusion. In February 2023, the KFTC imposed a total fine of 42.3 billion Won (approximately 33 million US dollars) on the four German automakers – Mercedes-Benz, BMW, Audi, and Volkswagen – for colluding on emission reduction technology applied to their diesel passenger cars. [^34] The KFTC’s ruling is available at https://www.ftc.go.kr/www/selectBbsNttView.do?pageUnit=10&pageIndex=92&searchCnd=all&key=12&bordCd=3&searchCtgry=01,02&nttSn=42758. Here, we investigate whether the German automakers also colluded on prices.
Drawing on province/year/product-level data on prices and attributes for 776 products from 13 brands between 2012 and 2023, [^35] We define a product as a unique combination of nameplate and fuel type. Appendix LABEL:app:_data_appendix provides details on the raw data and sample construction. we conduct pairwise hypothesis tests by running the following first-stage price regressions:
$p_{jt}=\gamma^{h}\cdot x_{jt}+\bm{\theta}^{h}\mathbf{z}_{jt}^{h}+\psi_{j}+\psi_{y}^{\textrm{fuel}}+\psi_{r}+e^{h}_{jt},$ (19)
where $x_{jt}$ represents four exogenous characteristics of product $j$ in market (year-province) $t$: fuel economy (km per 1,000 Won), acceleration (horsepower/weight), size (width×$\times$×length×$\times$×height), and a constant term. The regression model (19) also includes controls for product, year-fuel type (to account for differential trends in government policies regarding fuel types), and province, denoted as $\psi_{j}$, $\psi_{y}^{\textrm{fuel}}$, and $\psi_{r}$.
As for the exogenous instruments under conduct hypothesis $h\in\{H_{0},H_{1}\}$, denoted by $\mathbf{z}_{jt}^{h}$, we use the four exogenous product attributes to construct four summation IVs (sums over own‐firm products based on the effective firm index). In addition, we incorporate a two‐level nesting structure based on the car segment (small/compact, midsize, mid luxury, large/luxury, small SUV, standard SUV, large/luxury SUV) and fuel type (gasoline, diesel, LPG, EV, HEV), [^36] Note that while the nesting order is important for demand estimation, it is irrelevant to our testing procedure. and create 12 additional summation IVs: sums over own‐firm products in the same segment, sums over own‐firm products with the same fuel type, and sums over own‐firm products with the same fuel type in the same segment. [^37] We obtain testing results consistent with those reported in Table 4 when we proceed without any nesting structures and use just the four sums over own‐firm products as IVs.
We report the testing results (clustered by market) in the upper panel of Table 4 where a positive value greater than 1.65 (a negative value less than -1.65) constitutes evidence for the row (column) conduct model. The results indicate that, as expected, joint profit maximization by the parent company is more consistent with the data than own-profit maximization. Moreover, our test rejects all forms of price collusion against joint profit maximization by the parent company. These results, robust to alternative choices of fixed effects and clustering, align with the fact that no antitrust cases have been filed against the car manufacturers in South Korea in recent years, except for the non-price collusion case described above.
Table 4: Testing firm conduct in the two differentiated product markets
[…]
1. Brand ($F=13$) -3.718 1.013 0.335 4.940
2. Parent company ($F_{c}=9$) 3.718 5.404 2.229 7.104
3. German automakers ($F_{c}=7$) -1.013 -5.404 -0.343 4.954
4. Domestic/Foreign ($F_{c}=2$) -0.335 -2.229 0.343 4.168
5. Full collusion ($F_{c}=1$) -4.940 -7.104 -4.954 -4.168
[…]
1. Competition ($F=4$) 18.727 17.546 18.411 2.619
2. Full collusion ($F_{c}=1$) -18.727 -6.324 -5.444 -22.873
3. N-O-P ($F_{c}=2$) -17.546 6.324 3.747 18.316
4. N-O-S ($F_{c}=2$) -18.411 5.444 -3.747 21.808
5. N-P-S ($F_{c}=2$) -2.619 22.873 -18.316 -21.808
Notes: The table reports the results of pairwise hypothesis tests using our testing framework. Results for the new passenger car market are in the upper panel, while those for the instant noodles market are in the bottom panel. The names of the four ramen producers – Nongshim, Ottogi, Paldo, and Samyang – are abbreviated using their initial letters: N, O, P, and S. A positive value greater than 1.65 (a negative value less than -1.65) constitutes evidence for the row (column, respectively) conduct model.
8.2 Instant noodles market
The South Korean instant noodles (ramen) market is highly concentrated, with four firms – Nongshim, Ottogi, Samyang, and Paldo – accounting for over 90 percent of total sales. In March 2012, the KFTC fined them 135 billion won (approximately 120 million US dollars) for price collusion. In particular, Nongshim, the market leader with more than a 50% sales share, was fined 100 billion won for leading the collusion. However, this ruling was overturned by the Supreme Court of Korea in December 2015.383838The rulings of the KFTC and the Supreme Court are available at https://www.ftc.go.kr/www/selectBbsNttView.do?pageUnit=10&pageIndex=507&searchCnd=all&key=12&bordCd=3&searchCtgry=01,02&nttSn=37962 and https://www.scourt.go.kr/supreme/news/NewsViewAction2.work?seqnum=5081&gubun=4&searchOption=&searchWord=, respectively. Given that the KFTC remains suspicious of these firms and has been closely monitoring ramen prices, [^39] For example, the KFTC considered investigating collusion among the ramen manufacturers in 2023, citing that the ramen prices, which increased in the aftermath of the COVID-19 outbreak, did not decrease despite falling flour prices since late 2022: https://www.newsis.com/view/NISX20230623_0002349874. we test various collusion hypotheses, with Nongshim as the collusion leader, against the hypothesis of own-profit maximization. Specifically, we consider full collusion and three cases in which Nongshim coordinates prices with two other firms.
The data used for our test comprise region/year-month/product-level prices and product-level attributes of the 70 best-selling instant noodle products from 2010 to 2019.404040A region is composed of two or more adjacent provinces. See Appendix LABEL:app:_data_appendix for data details. For each of the seven product attributes – cholesterol, calorie, sugar, fat, protein, sodium contents, and a constant term – we calculate the sum over own-firm products (based on the effective firm index) and use them as exogenous instruments. In addition, given that the ramen market is differentiated along two dimensions (hong2023does; kim2024collusion), that is, package (pouch vs. cup) and soup type (red soup, white soup, soupless), [^41] We exclude cold noodle products from the analysis because their demand exhibits seasonality distinct from that of the products included in the analysis. we construct 21 additional IVs: sums over own‐firm products that share the same package, sums over own‐firm products that share the same soup type, and sums over own‐firm products that share both the same package and soup type. Then, we regress the price on product, time (year-month), and region dummy variables, along with the 28 IVs, for each conduct scenario. [^42] Product attributes are absorbed by the product dummies.
According to the testing results (clustered by market) presented in the bottom panel of Table 4, our test rejects all collusion hypotheses considered in favor of the own-profit maximization. [^43] We also consider other collusion scenarios in which Nongshim, along with another firm, coordinate prices, and obtain results in favor of price competition. These results align with the 2015 Supreme Court ruling, as well as the findings of kim2024collusion concluding that the observed markups are too low to support any collusive behavior.
9 Conclusion
Correctly assessing industry conduct is essential for establishing antitrust policy and evaluating market efficiency. Existing approaches to testing firm conduct often suffer from reduced testing power due to model misspecification and challenges in demand estimation. In this article, we propose a practical and powerful testing procedure that circumvents these limitations. Our method, built upon the Rivers and Vuong (RV) non-nested model selection framework, simply compares the performance of two BLP-style instrument sets – competition IVs and collusion IVs – in first-stage price regressions and interprets statistically significant results as evidence of either competitive or collusive behavior among firms.
Through extensive Monte Carlo simulations, we evaluate the finite-sample performance of our test statistic under various market conditions, characterized by different levels of market concentration, collusion, and internalization of colluding partners’ profits. The results show that our method is robust to model misspecification, alternative functional forms for instruments, and data limitations, performing comparably to, or better than, existing approaches in detecting collusion across various collusive scenarios. The simplicity of our approach despite its high testing power makes it a practical tool for the preliminary diagnosis of industry conduct. By complementing existing methods, our testing framework provides researchers and regulatory authorities with an efficient and effective way to assess firm behavior and guide antitrust interventions.
Moreover, our procedure can aid in designing more effective BLP-style instruments: once collusion is detected, researchers are advised to incorporate the product characteristics of colluding partners when building own-firm instruments while excluding them from other-firm instruments. This approach would strengthen the identification power of the instruments in demand estimation.